OpenStudy (anonymous):

HELP 1.Find the margin of error when p = 83% and n = 300. Round your answer to the nearest hundredth. 2.According to a recent survey, 78% of adults in the world are happy with their air quality. The margin of error was 0.03%. Approximately how many people were surveyed?

OpenStudy (superhelp101):

Well we know that the margin of error of a mean or standard error of a proportion is given by this : sqrt( var(x) / n), sometimes written as sd(x)/ sqrt(n) So the variance of a proportion is p ( 1 - p ) = pq = .83 x .17 = .14 The standard error is given by sqrt( .14 / 300 ) = .02

OpenStudy (superhelp101):

how it helps :D

OpenStudy (anonymous):

1) Margin of error = Confidence coefficient*Standard error of p Margin of error = 1.96*0.0217 = 0.04